# Deriving pulsar pair-production multiplicities from pulsar wind nebulae using H.E.S.S. and LHAASO observations

### v1

@@ 3 Results
For 26 pulsars in the Australia Telescope National Facility ATNF catalogue for which the region has been observed by H.E.S.S., we can place constraints on the lower limit of $\langle\kappa\rangle$. We do this by assuming that the birth period is in the range 10-50 ms, based on the reasoning that pulsars with birth periods below 10 ms are unlikely to exist (Perna et al., 2008), and PWNe associated with longer period pulsars are unlikely to produce gamma-ray emission. The resulting ranges for the possible lower-limit of $\langle\kappa\rangle$ are shown in Figure 1, with numerical values provided in Appendix A.
[…]
$E_{\mathrm{CR}}\approx 1.2\times 10^{20}A_{56}\eta\langle\kappa\rangle_{\rm lim,4}I_{45}B_{13}^{-1}R_{\star,6}^{-3}\tau_{7.5}^{-1}$ (11)
(Kotera et al., 2015), where $E_{\mathrm{CR}}$ is the iron nuclei energy which we assume to be $3\,\mathrm{PeV}$, roughly equivalent to the cosmic ray ‘knee’ (Blasi, 2013; Grenier et al., 2015), $A_{56}=1$ is the mass number of the stripped particles normalised to 56, and $B,R_{\star},I$ and $\tau$ are the magnetic field strength, radius of the pulsar, moment of inertia and age of the system respectively (with subscripts $x$ representing normalisation to $10^{x}$ in cgs units). $I_{45}$ and $R_{\star,6}$ we assume to be 1 as they are non-trivial to determine (Özel & Freire, 2016), $\tau_{7.5}$ we calculate using equation (7) and $B_{13}$ we obtain using the equation
$B=3.2\times 10^{19}\sqrt{P_{t}\dot{P_{t}}}\,\mathrm{G}$ (12)
(Gaensler & Slane, 2006). $\eta\leq 1$ in equation (11) is a luminosity conversion efficiency factor that is effectively unknown. Previous theoretical studies considering hadronic acceleration by pulsars have used values ranging from 0.1 to 1 (Kotera, 2014), and it could feasibly be of order 0.01 (de Oña Wilhelmi et al., 2022), although de Oña Wilhelmi et al. (2022) show that for many known gamma-ray sources it appears to be in the range 0.1 to 1 for particles downstream of the wind termination shock. As such, the conclusion reached for whether hadrons can escape into the pulsar wind for a given pulsar depends on the adopted value of $\eta$, as the intersection point between our derived $P_{0}$ value and the curve of $\langle\kappa\rangle(P_{0})$ can fall above or below the derived value for $\langle\kappa\rangle_{\rm lim}$.
The derived initial periods, pair multiplicities and values for $\langle\kappa\rangle_{\rm lim}$ for the six H.E.S.S. sources where we have both an estimate of $R_{PWN}/R_{SNR}$ and $N_{el}$ are shown in Figure 2 for assumed values for $\eta$ of 1 and 0.1. The calculated values for $P_{0}$ and $\langle\kappa\rangle$ for these sources is summarised in Table 3. We can confidently not exclude the possibility of hadronic escape into the wind for the pulsars J0835-4510 and J1930+1852, as they have intersection points between their $\langle\kappa\rangle(P_{0})$ curve and the modelled value of $P_{0}$ below the marker representing $\langle\kappa\rangle_{\rm lim}$ in the maximal case of assuming $\eta=1$. However, we also cannot completely exclude this scenario for the other four H.E.S.S. pulsars depending on the $\eta$ value assumed. That said, for J1747-2809 $\eta$ would be required to be order $10^{-3}$, so it is unlikely that hadrons escape into the wind for this source. The result that a higher efficiency factor for the production of hadrons at the pulsar corresponds to a greater likelihood of said hadrons escaping into the wind is expected, but our results quantify the effect of this for the first time. A previous claim in Kotera et al. (2015) that there is a rough value of $\langle\kappa\rangle_{\rm lim}$ of $2m_{p}/m_{e}\approx 3672$ for determining if hadrons can escape into the wind, for the specific hypothetical case of a young, $P_{0}\leq 10\,\mathrm{ms}$ pulsar producing ultra-high-energy cosmic rays, is probably therefore only correct to within 1-2 orders of magnitude. This is because, as can be seen from Figure 2, the constraints on hadronic escape reached for any particular source have that order uncertainty, depending on the measured or assumed values of $P$, $P_{0}$ and $\eta$. Despite differences in our modelling, we obtain order-of-magnitude similarity to the curve derived in de Jager (2007) for PSR B1509-58 (PSR J1513-5908), and whilst for the Vela Pulsar (PSR J0835-4510) our lower limit curve is within the permitted range predicted in that work, the precise level of agreement depends on scaling to the magnetic field strength in the region. However, some discrepancies for the results for these two pulsars considered in both our work and de Jager (2007) are to be expected, as we base our results on more recent observational data and PWN modelling, as referenced in Table 2.
With the exception of PSR J1833-1034, we obtain birth periods that are consistent within a factor two of previous estimates. The birth period we derive for PSR J1833-1034 of 33.0 ms is in stark contrast to the value of ¿55 ms derived by Camilo et al. (2006); this is likely due to Equation 1 not being suitable for a source that is potentially of age $<1$ kyr (much lower than its characteristic age of $\tau_{c}\approx 4.8$ kyr) as suggested by Camilo et al. (2009). The validity of our pair-production multiplicity for this source is therefore also questionable. The birth period for PSR J1513-5908 we derive (15.2 ms) is in good agreement with the value derived by Abdo et al. (2010) of 16 ms. Similarly for the Vela Pulsar PSR J0835-4510 the value of 10.9  ms we obtain is in reasonable agreement with the value from Helfand et al. (2001) of 6 ms. The birth period of 47.9 ms we obtain for PSR J1747-2809 is consistent with the prediction of Camilo et al. (2009) who proposed it to be ¿40 ms. It has been postulated that pulsars with approximately millisecond birth periods could explain the ultra-high energy all-particle cosmic ray flux if they exist (Kotera et al., 2015); we find no pulsars with a $\sim\mathrm{ms}$ birth period in our sample. This is consistent with some previous studies utilising archival X-ray data (Perna et al., 2008).
The results of the modelling of the four LHAASO sources can be seen in Figure 3. We cannot attempt to derive values for $\langle\kappa\rangle_{\rm lim}$ as we do not have radio observations for these targets. But if we assume the potential range of $\langle\kappa\rangle_{\rm lim}$ values for these systems is comparable to those we have determined with the H.E.S.S. data, we cannot exclude the possibility of hadronic particles reaching the pulsar wind for any of these systems either.
@@
Figure 2: The intersections of our estimates for $P_{0}$ using Equation 2, and the curves of $\langle\kappa\rangle(P_{0})$ as estimated using Equations 3, 6 and 7 (the latter is constrained such that $P_{0}<P_{t}$) are shown by circular markers. The intersection location represents the lower limit for the pair-production multiplicity, see Table 3 for the associated values. Derived values for the pair-production multiplicity limit $\langle\kappa\rangle_{\rm lim}$ for each pulsar are shown as lower limit markers in matching colours to the multiplicity curve. The top panel assumes a luminosity conversion efficiency of $\eta=1$ in the determination of $\langle\kappa\rangle_{\rm lim}$, whereas the lower panel assumes $\eta=0.1$.
J1833-1034 33.0 $33.9$ $45.6$ $51.8$ 1476
J1513-5908 15.2 $17.0$ $5.14$ $8.35$ 809
J1930+1852 41.3 $12.0$ $5.06$ $11.0$ 432
J1846-0258 50.8 $8.13$ $1.62$ $1.87$ 1061
J0835-4510 10.9 $6.92$ $18.7$ $1.87$ 174
J1747-2809 47.9 $42.7$ $125$ $71.4$ 29328
[…]
Figure 3: The calculated lower limits on the pair-production multiplicity for the four pulsars co-incident with LHAASO sources.
4 Discussion
4.1 Implications for production of cosmic rays
We cannot safely exclude the possibility of hadrons escaping the pulsar surface into the wind for any of our considered sources. Though it should be noted that the values of $\langle\kappa\rangle$ we observe for all sources are within likely theoretical constraints (given $P_{0}$ is not likely to be $\gg$50ms), which cap the maximum pair production multiplicity to be a few $10^{5}$ (Timokhin & Harding, 2019). The $\langle\kappa\rangle$ curve we observe for PSR J1849-0001 is significantly below unity, but similar findings were encountered by de Jager (2007), who found that the Vela Pulsar (PSR J0835-4510) had a pair production multiplicity less than unity (depending on the assumptions made).
4.2 The phase-space of free parameters for the H.E.S.S. sources
As our work is based on the previous findings of Giacinti et al. (2020), which did not provide estimates of the uncertainties on the PWN and SNR radii, we cannot perform a full, robust uncertainty analysis. Instead, to investigate the effect of various assumptions in our models, we examine the effect of changing parameters in our model for the two pulsars PSR J1747-2809 and PSR J1846-0258 (to probe the minimum and maximum values of $P_{t}$ respectively). The results of this are shown in Figures 4 and 5.
The selected value of $I$ appears to have an approximately order of magnitude effect on the point of intersection between the pair production multiplicity curve and the pulsar birth period estimate for the short period pulsar PSR J1747-2809 seen in Figure 4, but very little influence upon the derived pair-production multiplicity as seen in Figure 5 for PSR J1846-0258. Variations in the moment of inertia $I$ correspond to differences in the internal structure, and can be used to constrain the equation of state for dense matter Ravenhall & Pethick (1994). The results for both pulsars suggest there is a roughly factor 2 impact on the derived $P_{0}$ within the band of reasonable potential values. Changing the value of $E_{2}$ is seen to only have a modest effect on the value of $\langle\kappa\rangle$ in Figures 4 and 5. Further increasing $E_{2}$ would lead to a resulting decrease in $\langle\kappa\rangle$. Choosing 10 TeV for this value appears to be reasonable as a lower limit given our relative agreement for the Vela pulsar and PSR B1509-58 with de Jager (2007), where the pair-production multiplicity is calculated differently. Changing the value of $\Gamma$ by 0.4 appears to have an approximately factor two effect on the derived $\langle\kappa\rangle$. The true pair production multiplicity value is very likely to be within this range, as a value of 2.2 is expected for particles at an ultra-relativistic shock undergoing isotropic scattering (Achterberg et al., 2001). A similar magnitude effect is seen when varying $n$ for both pulsars, this likely dominates the uncertainty for shorter birth period pulsars. But as this braking index has only been measured for a very small population of pulsars, by convention we adopt a value of 3 as has been adopted by much of the literature (Gaensler & Slane, 2006).
4.3 Prospects for future observations
The large beaming fraction (up to 0.92) for gamma-rays from high $\dot{E}_{t}$ pulsars (Johnston et al., 2020) suggests that observations of gamma-ray pulsars are nearly complete. This implies conclusions of this study regarding the contribution of pulsars to the hadronic cosmic ray spectra are unlikely to change with the advent of future gamma-ray instruments such as the Cherenkov Telescope Array (Cherenkov Telescope Array Consortium et al., 2019) or the Southern Wide-Field Gamma-Ray Observatory (Albert et al., 2019). That said, state-of-the-art and future radio observations by observatories such as MeerKAT (Goedhart et al., 2023) and the upcoming Square Kilometer Array (Carilli & Rawlings, 2004) could allow for more PWN and SNR radii, from systems with lower surface brightness and larger extent, to be measured. This means the pair production multiplicities and birth period could potentially be derived for more pulsars. This, in combination with multi-wavelength modelling, would allow for more stringent constraints on hadronic contributions to the gamma-ray emission from PWNe.
Figure 4: The effect of varying $I$ (top), $E_{2}$ (upper middle), $\Gamma$ (lower middle) and $n$ (bottom) for PSR J1747-2809, which has the lowest $P_{t}$ (47.9 ms) in our sample of pulsars with a constrained $P_{0}$.
Figure 5: The effect of varying $I$ (top), $E_{2}$ (upper middle), $\Gamma$ (lower middle) and $n$ (bottom) for PSR J1846-0258, which has the highest $P_{t}$ (326.6 ms) in our sample of pulsars with a constrained $P_{0}$.
5 Conclusions
In this study we have placed constraints on the pair production multiplicity for 26 pulsars observed by H.E.S.S., deriving lower limits for the average pair production multiplicities and birth periods for six of these pulsars. We also placed lower limits on the pair production multiplicities for four pulsars co-incident with sources observed by LHAASO based on their very-high-energy gamma-ray emission. We find that the possibility of hadrons escaping the pulsar into the wind cannot be conclusively excluded for any source we consider. Nevertheless, for J1747-2809 the pulsar for which we infer the highest pair-production multiplicity, the escape of hadrons into the wind is essentially ruled out for values of $\eta\geq 0.1$. This implies that hadrons are only released into the wind of J1747-2809 if $\eta\leq 0.1$, which is too low for any hadronic component to be energetically significant.
Acknowledgements.
Appendix A Numerical estimates of $\langle\kappa\rangle$
Numerical bounds for the pair-production multiplicities for the 26 pulsars in our combined H.E.S.S./ATNF sample are shown in Table 3.
@@ Appendix A Numerical estimates of $\langle\kappa\rangle$
Table 3: Estimates of the maximum and minimum value of $\langle\kappa\rangle$ for the H.E.S.S./ATNF sample.

### v3

@@ 3 Results
For 26 pulsars in the Australia Telescope National Facility ATNF catalogue, whose regions have been observed by H.E.S.S., we can place constraints on the lower limit of $\langle\kappa\rangle$. We do this by assuming that the birth period is in the range 10-50 ms, based on the reasoning that pulsars with birth periods below 10 ms are unlikely to exist (Perna et al., 2008) and PWNe associated with longer period pulsars are unlikely to produce gamma-ray emission. The resulting ranges for the possible lower-limit of $\langle\kappa\rangle$ are shown in Fig. 1, with numerical values provided in Table 2.
@@
Table 2: Estimates of the maximum and minimum values of the lower limit for $\langle\kappa\rangle$ for the H.E.S.S./ATNF sample.
@@
[…]
$E_{\mathrm{CR}}\approx 1.2\times 10^{20}A_{56}\eta\langle\kappa\rangle_{\rm lim,4}I_{45}B_{13}^{-1}R_{\star,6}^{-3}\tau_{7.5}^{-1}\,\mathrm{eV},$ (11)
where $E_{\mathrm{CR}}$ is the iron nuclei energy which we assume to be $3\,\mathrm{PeV}$, roughly equivalent to the cosmic ray ‘knee’ (Blasi, 2013; Grenier et al., 2015), $A_{56}=1$ is the mass number of the stripped particles normalised to 56, and $B,R_{\star},I$, and $\tau$ are the magnetic field strength, radius of the pulsar, moment of inertia, and age of the system, respectively (with subscripts $x$ representing normalisation to $10^{x}$ in cgs units). $I_{45}$ and $R_{\star,6}$ we assumed to be equal to 1, as they are non-trivial to determine (Özel & Freire, 2016), whereas we calculated $\tau_{7.5}$ using Eq. (7) and $B_{13}$ was obtained using the equation:
$B=3.2\times 10^{19}\sqrt{P_{t}\dot{P_{t}}}\,\mathrm{G}$ (12)
(Gaensler & Slane, 2006). Here, $\eta\leq 1$ in Eq. (11) is a luminosity conversion efficiency factor that is effectively unknown. Previous theoretical studies considering hadronic acceleration by pulsars have used values ranging from 0.1 to 1 (Kotera, 2014) and it could feasibly be of the order of 0.01 (de Oña Wilhelmi et al., 2022), although de Oña Wilhelmi et al. (2022) showed that for many known gamma-ray sources, it appears to be in the range 0.1 to 1 for particles downstream of the wind termination shock. As such, the conclusion reached for whether hadrons can escape into the pulsar wind for a given pulsar depends on the adopted value of $\eta$, as the intersection point between our derived $P_{0}$ value and the curve of $\langle\kappa\rangle(P_{0})$ can fall above or below the derived value for $\langle\kappa\rangle_{\rm lim}$.
The derived initial periods, pair multiplicities and values for $\langle\kappa\rangle_{\rm lim}$ for the six H.E.S.S. sources, where we have both an estimate of $R_{PWN}/R_{SNR}$ and $N_{el}$ are shown in Fig. 2 for assumed values for $\eta$ of 1 and 0.1. The calculated values for $P_{0}$ and $\langle\kappa\rangle$ for these sources is summarised in Table 4. We can confidently not exclude the possibility of hadronic escape into the wind for the pulsars J0835-4510 and J1930+1852, as they have intersection points between their $\langle\kappa\rangle(P_{0})$ curve and the modelled value of $P_{0}$ below the marker representing $\langle\kappa\rangle_{\rm lim}$ in the maximal case of assuming $\eta=1$. However, we also cannot completely exclude this scenario for the other four H.E.S.S. pulsars depending on the $\eta$ value assumed. That said, for J1747-2809, the value of $\eta$ would have to be of the order of $10^{-3}$, so it is unlikely that hadrons escape into the wind for this source. The result that a higher efficiency factor for the production of hadrons at the pulsar corresponds to a greater likelihood of said hadrons escaping into the wind is expected, but our results quantify the effect of this for the first time. It was previously claimed in Kotera et al. (2015) that there is a rough value of $\langle\kappa\rangle_{\rm lim}$ of $2m_{p}/m_{e}\approx 3672$ for determining whether hadrons can escape into the wind, based on the specific hypothetical case of a young, $P_{0}\leq 10\,\mathrm{ms}$ pulsar producing ultra-high-energy cosmic rays. This claim is is only likely to be correct to within one to two orders of magnitude. This is because (as seen from Fig. 2) the constraints on hadronic escape reached for any particular source have that order uncertainty, depending on the measured or assumed values of $P$, $P_{0}$, and $\eta$. Despite the differences in our modelling, we obtained an order-of-magnitude similarity to the curve derived in de Jager (2007) for PSR B1509-58 (PSR J1513-5908). For the Vela Pulsar (PSR J0835-4510), our lower limit curve is within the permitted range predicted in that work, however, the precise level of agreement depends on scaling to the magnetic field strength in the region. However, some discrepancies for the results for these two pulsars considered in both our work and de Jager (2007) are to be expected, as we have based our results on more recent observational data and PWN modelling, as referenced in Table 2.
With the exception of PSR J1833-1034, we obtained birth periods that are consistent within a factor two of previous estimates. The birth period we derived for PSR J1833-1034 of 33.0 ms is in stark contrast to the value of ¿55 ms derived by Camilo et al. (2006); this is likely due to Eq. 1 not being suitable for a source that is potentially of age $<1$ kyr (much lower than its characteristic age of $\tau_{c}\approx 4.8$ kyr), as suggested by Camilo et al. (2009). The validity of our pair-production multiplicity for this source is therefore also questionable. The birth period for PSR J1513-5908 we derive (15.2 ms) is in good agreement with the value derived by Abdo et al. (2010) of 16 ms. Similarly for the Vela Pulsar PSR J0835-4510 the value of 10.9  ms we obtain is in reasonable agreement with the value from Helfand et al. (2001) of 6 ms. The birth period of 47.9 ms we obtain for PSR J1747-2809 is consistent with the prediction of Camilo et al. (2009) who proposed it to be ¿40 ms. It has been postulated that pulsars with approximately millisecond birth periods could explain the ultra-high energy all-particle cosmic ray flux if they exist (Kotera et al., 2015); we find no pulsars with a $\sim\mathrm{ms}$ birth period in our sample. This is consistent with some previous studies utilising archival X-ray data (Perna et al., 2008).
The results of the modelling of the four LHAASO sources can be seen in Fig. 3. We cannot attempt to derive values for $\langle\kappa\rangle_{\rm lim}$ as we do not have radio observations for these targets. However, if we assume the potential range of $\langle\kappa\rangle_{\rm lim}$ values for these systems is comparable to those we have determined with the H.E.S.S. data, we cannot exclude the possibility of hadronic particles reaching the pulsar wind for any of these systems either.
Figure 2: Intersections of our estimates for $P_{0}$ using Eq. 2, and the curves of $\langle\kappa\rangle(P_{0})$ as estimated using Eqs. 3, 6, and 7 (the latter is constrained such that $P_{0}<P_{t}$) are shown by circular markers. The intersection location represents the lower limit for the pair-production multiplicity (see Table 4) for the associated values. Derived values for the pair-production multiplicity limit $\langle\kappa\rangle_{\rm lim}$ for each pulsar are shown as lower limit markers in matching colours to the multiplicity curve. The top panel assumes a luminosity conversion efficiency of $\eta=1$ in the determination of $\langle\kappa\rangle_{\rm lim}$, whereas the lower panel assumes $\eta=0.1$.
J1833-1034 33.0 $33.9$ $45.6$ $51.8$ 1476
J1513-5908 15.2 $17.0$ $5.14$ $8.35$ 809
J1930+1852 41.3 $12.0$ $5.06$ $11.0$ 432
J1846-0258 50.8 $8.13$ $1.62$ $1.87$ 1061
J0835-4510 10.9 $6.92$ $18.7$ $24.7$ 174
J1747-2809 47.9 $42.7$ $125$ $71.4$ 29328
[…]
Figure 3: Calculated lower limits on the pair-production multiplicity for the four pulsars co-incident with LHAASO sources.
4 Discussion
4.1 Implications for production of cosmic rays
We cannot safely exclude the possibility of hadrons escaping the pulsar surface into the wind for any of our considered sources. However, it should be noted that the values of $\langle\kappa\rangle$ we observe for all sources are within likely theoretical constraints (given $P_{0}$ is not likely to be $\gg$50ms), which cap the maximum pair production multiplicity to be a few times $10^{5}$ (Timokhin & Harding, 2019). The $\langle\kappa\rangle$ curve we observe for PSR J1849-0001 is significantly below unity, but similar findings were encountered by de Jager (2007), who found that the Vela Pulsar (PSR J0835-4510) had a pair production multiplicity less than unity (depending on the assumptions made).
4.2 The phase-space of free parameters for the H.E.S.S. sources
As our work is based on the previous findings of Giacinti et al. (2020), which did not provide estimates of the uncertainties on the PWN and SNR radii, we cannot perform a full, robust uncertainty analysis. Instead, to investigate the effect of various assumptions in our models, we have examined the effect of changing parameters in our model for the two pulsars PSR J1747-2809 and PSR J1846-0258 (to probe the minimum and maximum values of $P_{t}$ respectively). The results of this are shown in Figs. 4 and 5.
The selected value of $I$ appears to have an approximately order of magnitude effect on the point of intersection between the pair production multiplicity curve and the pulsar birth period estimate for the short period pulsar PSR J1747-2809 (seen in Fig. 4); however, it exerts very little influence upon the derived pair-production multiplicity, as seen in Fig. 5 for PSR J1846-0258. Variations in the moment of inertia $I$ correspond to differences in the internal structure and can be used to constrain the equation of state for dense matter Ravenhall & Pethick (1994). The results for both pulsars suggest there is an impact on the derived $P_{0}$ by roughly a factor of 2, within the band of reasonable potential values. Changing the value of $E_{2}$ is seen to only have a modest effect on the value of $\langle\kappa\rangle$ in Figs. 4 and 5. Further increasing $E_{2}$ would lead to a resulting decrease in $\langle\kappa\rangle$. Choosing 10 TeV for this value appears to be reasonable as a lower limit given our relative agreement for the Vela pulsar and PSR B1509-58 with de Jager (2007), where the pair-production multiplicity is calculated differently. Changing the value of $\Gamma$ by 0.4 appears to have an approximately factor two effect on the derived $\langle\kappa\rangle$. The true pair production multiplicity value is very likely to be within this range, as a value of 2.2 is expected for particles at an ultra-relativistic shock undergoing isotropic scattering (Achterberg et al., 2001). A similar magnitude effect is seen when varying $n$ for both pulsars and this is likely to dominate the uncertainty for shorter birth period pulsars. However, as this braking index has only been measured for a very small population of pulsars, by convention, we adopted a value of 3, previously adopted by many works in the literature as well (Gaensler & Slane, 2006).
4.3 Prospects for future observations
The large beaming fraction (up to 0.92) for gamma-rays from high $\dot{E}_{t}$ pulsars (Johnston et al., 2020) suggests that observations of gamma-ray pulsars are nearly complete. This implies conclusions of this study regarding the contribution of pulsars to the hadronic cosmic ray spectra are unlikely to change with the advent of future gamma-ray instruments, such as the Cherenkov Telescope Array (Cherenkov Telescope Array Consortium et al., 2019) or the Southern Wide-Field Gamma-Ray Observatory (Albert et al., 2019). That said, state-of-the-art and future radio observations by observatories such as MeerKAT (Goedhart et al., 2023) and the upcoming Square Kilometer Array (Carilli & Rawlings, 2004) could allow for more PWN and SNR radii, from systems with lower surface brightness and larger extent, that are yet to be measured. This means the pair production multiplicities and birth period could potentially be derived for more pulsars. In combination with multi-wavelength modelling, this would allow for more stringent constraints on hadronic contributions to the gamma-ray emission from PWNe.
Figure 4: Effect of varying $I$ (top), $E_{2}$ (upper middle), $\Gamma$ (lower middle), and $n$ (bottom) for PSR J1747-2809, which has the lowest $P_{t}$ (47.9 ms) in our sample of pulsars with a constrained $P_{0}$.
Figure 5: Effect of varying $I$ (top), $E_{2}$ (upper middle), $\Gamma$ (lower middle) and $n$ (bottom) for PSR J1846-0258, which has the highest $P_{t}$ (326.6 ms) in our sample of pulsars with a constrained $P_{0}$.
5 Conclusions
In this study, we have placed constraints on the pair production multiplicity for 26 pulsars observed by H.E.S.S., deriving lower limits for the average pair production multiplicities and birth periods for six of these pulsars. We also set lower limits on the pair production multiplicities for four pulsars that are co-incident with sources observed by LHAASO, based on their very-high-energy gamma-ray emission. We find that the possibility of hadrons escaping the pulsar into the wind cannot be conclusively excluded for any source we consider here. Nevertheless, for J1747-2809 (the pulsar for which we infer the highest pair-production multiplicity), the escape of hadrons into the wind is essentially ruled out for values of $\eta\geq 0.1$. This implies that hadrons are only released into the wind of J1747-2809 if $\eta\leq 0.1$, which is too low for any hadronic component to be energetically significant.
Acknowledgements.